OpenStudy (anonymous):

Write an equation in point-slope form for the line through the given point with the given slope. (8, 3); m = 6 (1 point)

jimthompson5910 (jim_thompson5910):

y - y1 = m(x-x1) y - 3 = 6(x-8)

OpenStudy (anonymous):

y=6x-45